The length of a rectangle exceeds it width by 17 inches and the area 18 square inches. What are the length and width of the of the rectangle
Answer by stanbon(75887)

The length of a rectangle exceeds it width by 17 inches and the area 18 square inches. What are the length and width of the of the rectangle
--------------------
Let width be "x"
Then length is "x+17"
-----
Equation:
x(x+17) = 18
x^2 + 17x - 18 = 0
----
Factor:
(x+18)(x-1) = 0
---
Positive solution:
x = 1 inch (width)
x+17 = 18 inches (length)
=============================
Cheers,
Stan H.
===============